# Hybrid Type-II and Type-III seesaw model for the muon $g-2$ anomaly

### v1

@@ Abstract
In this work, we investigate the muon anomalous dipole moment $a_{\mu}$ in a model that extends the Standard Model with a scalar triplet and a vector-like lepton triplet. Different from previous studies, we find that there is still viable parameter space in this model to explain the discrepancy $\Delta a_{\mu}=a_{\mu}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})$. While being consistent with the current data of neutrino mass, precision electroweak measurements and the perturbativity of couplings, our model can provide new physics contribution $a_{\mu}^{\textrm{NP}}$ to cover the central region of $\Delta a_{\mu}$ with new scalar and fermion mass as low as around TeV. This mass scale is allowed by the current collider searches for doubly charged scalars and vector-like leptons, and they can be tested at future high energy and/or high luminosity colliders.
@@ I Introduction
The anomalous magnetic moment (AMM) of muon, denoted as $a_{\mu}\equiv(g-2)_{\mu}/2$, has been theoretically predicted in the Standard Model (SM) and experimentally measured both to very high precision. Since the $g-2$ experiment E821 performed at Brookhaven National Laboratory (BNL) released its data about two decades ago [1], a discrepancy $\Delta a_{\mu}=a_{\mu}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})$ has been existing and triggered rich phenomenological studies (see review [2, 3] and and references therein). Recently, on the theoretical side, a comprehensive summary of the most accurate SM prediction for $a_{\mu}$ is provided in [4] where the value is reported as $a_{\mu}(\mathrm{SM})=116591810(43)\times 10^{-11}$. On the experimental side, the Muon $g-2$ Experiment at Fermilab released its Run-1 dataset in 2021 and the experimental result was $a_{\mu}^{2021}(\mathrm{Exp})=116592061(41)\times 10^{-11}$ [5] when combining the data from Fermilab Run-1 and BNL. This resulted in a deviation from the SM prediction of $a_{\mu}^{2021}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})=(251\pm 59)\times 10^{-11}$ with a significance of $4.2\,\sigma$. In 2023, Fermilab released data from Run-2 and Run-3 leading to a new result after being combined with Run-1 and the BNL data as $a_{\mu}^{2023}(\mathrm{Exp})=116592059(22)\times 10^{-11}$ [5]. In this case the deviation from the SM prediction $\Delta a_{\mu}=a_{\mu}^{2023}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})=(249\pm 48)\times 10^{-11}$ achieved a significance of $5.1\,\sigma$. These results inspired phenomenological investigations in various new physics models [6, 7, 8, 9]  [^1] Recently, there have been some discussions on the discrepancies in the calculation of the hadronic vacuum polarization contribution to $a_{\mu}(\mathrm{SM})$ between the lattice QCD calculations and experimental data in measurements of $e^{-}e^{+}\to\pi^{-}\pi^{+}$, which seems to reduce $\Delta a_{\mu}$ to some extend [10, 11, 12]. .
In the framework of simplified models, contributions to $\Delta a_{\mu}$ can be generated at the 1-loop level by introducing several new physics fields with various spins and quantum numbers under the electroweak gauge group ${\rm SU(2)_{L}\times U(1)_{Y}}$. For discussions involving one, two, or three new fields, see the recent reviews [13, 14, 15, 16] and the references therein. The chiral structure of the $a_{\mu}$ can be described by the following effective tensor operator
$\delta\mathcal{L}^{\rm AMM}_{\rm eff}=-\frac{e_{\ell}a_{\ell}}{4m_{\ell}}\,\overline{\ell_{L}}\,\sigma^{\mu\nu}\,\ell_{R}\,F_{\mu\nu}\,+{\rm h.c.}\,,$ (1)
@@
in which the SM charged lepton $\ell$ should be understood as muon flavor. We can see that both the left-handed (LH) $\ell_{L}$ and the right-handed (RH) $\ell_{R}$ of muon are involved. More specifically, if one considers the new physical contribution to $a_{\mu}$ at 1-loop level, proper chiral flip is needed along the fermion line. This chiral flip can take place either in the external muon line realized by the SM Yukawa interaction outside the loop or in the internal fermion line inside the loop [15].
To explain $\Delta a_{\mu}$ it generally requires either new fields with light masses to avoid too much loop suppression or heavy particles inside the loop to generate chiral enhancement to the amplitude. However, light new fields are often constrained by electroweak precision observables, flavor physics, and collider searches, typical examples of which include the two-Higgs-doublet model [17] and the dark photon model [18]. Given the aformentioned experimental constraints many models are viable only in the parameter space where the new particles are heavy. Therefore, proper chiral enhancement is typically necessary to adequately explain $\Delta a_{\mu}$.
As pointed out in recent reviews [13, 16], many models fail to explain $\Delta a_{\mu}$ either due to the absence of chiral enhancement or the wrong sign of $\Delta a_{\mu}$ in the prediction (e.g. minimal seesaw models [19, 20]). Chiral enhancement can be achieved, for instance, through a chiral flip occurring on a fermionic field within the loop as manifested in supersymmetric models [21, 22]. These fermionic fields can also appear as charm/top quark in SM or vector-like quarks in leptoquark-extended models [23, 24, 25], or more generally as new physics fermions being Dirac, Majorana, or Weyl nature.
@@
It has been found in recent studies [14, 15, 16] that simplified models introducing two new fields with different spins to the SM are unable to generate the necessary chiral enhancement within the 1-loop structure of $a_{\mu}$. As a result, these models tend to confine the new particles to a low and compressed mass region in order to explain $\Delta a_{\mu}$ while avoiding collider constraints. However, an important consideration has been overlooked in the aforementioned analyses: the Yukawa interaction involving the SM Higgs doublet and the newly introduced fermion. The Yukawa interaction, a four-dimensional operator, naturally exists and can induce mass mixing between fermions in the SM and the new physics sector.
@@
In this work, we demonstrate that by incorporating the SM Higgs Yukawa interaction into the analysis, simplified models introducing two new fields to the SM can effectively explain $\Delta a_{\mu}$. As a concrete example, we consider a simplified model that extends the SM with a scalar triplet $(3,-1)$ and a vector-like lepton (VLL) triplet $(3,0)$ under the electroweak gauge group ${\rm SU(2)_{L}\times U(1)_{Y}}$. These two particles have been scrutinized in neutrino mass generation mechanisms as Type-II [26, 27, 28, 29, 30, 31] and Type-III [32, 33] seesaw models, respectively. However, in this work concentrating on the physics of $(g-2)_{\mu}$, we do not require our model to produce the experimentally suggested texture of neutrino mass matrix, but instead ensure the theoretically predicted neutrino mass remains to be negligibly small and close to zero.
@@
This paper is organized as follows. In Sec. II we articulate our model setup and the related parameters. In Sec. III we present our analytical and numerical results and we draw our conclusion in Sec. IV.
@@ II.1 Model setup
In our model we extend the SM with a scalar triplet $S$ and a vector-like lepton (VLL) triplet $F$, of which the representation $(n,\,Y)$ under the SM electroweak gauge group ${\rm\mathrm{SU(2)_{L}}\times\mathrm{U(1)_{Y}}}$ are
[…]
@@
$S\equiv\Big{[}\begin{array}[]{cc}\delta^{-}/\sqrt{2}&(\delta^{0}+v_{\delta})/\sqrt{2}\\ \delta^{--}&-\delta^{-}/\sqrt{2}\end{array}\Big{]}\,,\qquad F_{L}\equiv\Big{[}\begin{array}[]{cc}F_{L}^{0}/\sqrt{2}&F_{L}^{+}\\ F_{L}^{-}&-F_{L}^{0}/\sqrt{2}\end{array}\Big{]}\,.$ (7)
We consider the following mass and Yukawa terms in the Lagrangian which are most relevant to the physics of $(g-2)_{\mu}$
@@
in which $H$ is the SM scalar doublet and $L^{i,j}_{L}$ ($\ell^{i,j}_{R}$) is the LH (RH) lepton doublet (singlet) in SM with $i,j=1,2,3$ denoting the generation index. $F^{C}$ denotes the charge conjugation of the field $F$ satisfying $F^{C}\equiv C\overline{F}^{T}$ with $C\equiv i\gamma^{2}\gamma^{0}$. Note that in our notation the $C$ operator also acts on the representation space of SM gauge group.
After performing the expansion $H=[0,({v_{h}}+h)/\sqrt{2}]^{T}$ we can derive the following neutrino and charged lepton mass matrices
@@
Then, the Yukawa interactions can be written as
@@
For simplicity, we consider the scenario that only the second generation $i,j\equiv 2$ exist in Eq. (8), i.e. the mass mixing in Eq. (9) and Eq. (10) would generate mass eigenstates of physical charged lepton and neutrino for muon flavor, as well as heavy neutral and charged leptons in the new physics sector. Therefore, the indices $i,j$ will be dropped in the following. We can diagonalize the previous mass matrices through the following rotations
@@
In the above, the $s_{L,R}^{\nu,\,\ell}$ and $c_{L,R}^{\nu,\,\ell}$ are abbreviations of $\sin\theta_{L,R}^{\nu,\,\ell}$ and $\cos\theta_{L,R}^{\nu,\,\ell}$ when applicable. After the transformations in Eq. (11) we obtain the following Yukawa interactions in terms of mass eigenstates
@@
In general, there are interactions in the scalar sector involving $H$ and $S$ [34] as follows
$V(H,S)=$ $-m_{H}^{2}H^{{\dagger}}H+\frac{\lambda}{4}(H^{{\dagger}}H)^{2}+M_{S}^{2}\mathrm{Tr}[S^{{\dagger}}S]+[\mu_{HS}(H^{T}i\sigma_{2}SH)+\mathrm{h.c.}]$
@@
which would generate the mixing between $h$ and $\delta^{0}$ and result in the physical mass eigenstates including the SM Higgs scalar and heavy scalars in the new physics sector. Similar to Eq. (11) we can diagnolize the scalar mass matrix by performing the following rotation
@@
in which $s^{h},c^{h}$ are abbreviations of $\sin\theta^{h},\cos\theta^{h}$ following the similar convention as Eq. (11). The physical masses of $h,\delta^{0},\delta^{-},\delta^{--}$ will be labelled as $m_{h},m_{\delta^{0}},m_{\delta^{-}},m_{\delta^{--}}$ in later analysis.
@@ II.2 Relations of parameters
In Eq. (II.2) and Eq. (II.2), there are different equalities on $z_{L}v_{h}$ and $M_{F}$. Therefore, we can have the following identities
@@
  • $s_{R}^{\ell}$ characterizing the mixing of RH component of SM charged lepton $\ell$ with $F^{-}$ can be solved from Eq. (22) and Eq. (II.2) and turns out to be the following in the parameter regions chosen for our numerical calculation [^2] Note that in other models the suppression of $s^{\ell}_{R}$ compared to $s^{\ell}_{L}$ can be different from our model. Taking Type-III seesaw model as an example, the suppression is $\sim m_{\ell}/M_{F}$ [37, 38].
@@
    which means that the small angle approximation also holds for $\theta_{R}$ (see Eq. (II.3) and Eq. (II.3) later for more details). Therefore, in our model we have
@@
  • As the mass gap $\Delta m_{F}\equiv m_{F^{\pm}}-m_{F^{0}}$ in Eq. (24) in our model is negligibly small, $\Delta m_{F}$ does not alter the main decay signals of heavy fermions $F^{\pm},F^{0}$ as discussed in [39, 40]. Therefore, we can simply impose the similar lower bound of VLL mass to be
@@
II.3 Independent input parameters
Now we determine the physically reasonable choice of independent input parameters of our model.
In the scalar sector shown in Eq. (II.1), despite the rich parameters and phenomenology about $V(H,S)$ (see e.g. [41] for collider signal searches), we concentrate on the following input parameters which are the most relevant quantities to $(g-2)_{\mu}$
$\{\,s^{h}\,,\quad v_{\delta}\,,\quad m_{\delta}\,\}\,.$ (30)
Our requirements include the following ones.
  • $v_{h},v_{\delta}$ are required to satisfy [42, 43]
    $v\approx\sqrt{v_{h}^{2}+2v_{\delta}^{2}}\approx 246\,{\rm GeV}\,,$ (31)
    to generate a SM-like Higgs boson as discovered at the Large Hadron Collider (LHC) [44, 45].
  • To avoid the constraints $v_{\delta}\lesssim 5\,{\rm GeV}$ from the electroweak precision observables [46], in our numerical calculation we choose
    $v_{\delta}=5\,{\rm GeV}\,.$ (32)
  • Under the previous two conditions which leads to $v_{\delta}\ll v_{h}$, we can utilize the approximations between the model parameter $M_{S}$ in Eq. (II.1) and the masses of physical scalars $\delta^{0},\delta^{-},\delta^{--}$, i.e. $M_{S}\approx m_{\delta^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}$ [34, 41]. Hereafter we would use $m_{\delta}$ as the notation to denote
    $m_{\delta}\equiv m_{\delta^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}\,.$ (33)
    Considering that the current doubly charged scalar search sets a lower limit of mass to be around $1\,{\rm TeV}$ assuming decaying to SM leptons [47, 48], we take the following benchmark throughout this work

### v2

@@ Abstract
In this work we investigate the muon anomalous dipole moment $a_{\mu}$ in a model that extends the Standard Model with a scalar triplet and a lepton triplet. Different from previous studies, we find that there is still viable parameter space in this model to explain the discrepancy $\Delta a_{\mu}=a_{\mu}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})$. While being consistent with the current data of neutrino mass, electroweak precision measurements and the perturbativity of couplings, our model can provide new physics contribution $a_{\mu}^{\textrm{NP}}$ to cover the central region of $\Delta a_{\mu}$ with new scalar and lepton mass as low as around TeV. This mass scale is allowed by the current collider searches for doubly charged scalars and the lepton triplet, and they can be tested at future high energy and/or high luminosity colliders.
preprint: APCTP-Pre2024-017
@@ I Introduction
The anomalous magnetic moment (AMM) of muon, denoted as $a_{\mu}\equiv(g-2)_{\mu}/2$, has been theoretically predicted in the Standard Model (SM) and experimentally measured both to very high precision. Since the $g-2$ experiment E821 performed at Brookhaven National Laboratory (BNL) released its data about two decades ago [1], a discrepancy $\Delta a_{\mu}=a_{\mu}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})$ has been existing and triggered rich phenomenological studies (see reviews [2, 3] and and references therein). Recently, on the theoretical side, a comprehensive summary of the most accurate SM prediction for $a_{\mu}$ is provided in [4] where the value is reported as $a_{\mu}(\mathrm{SM})=116591810(43)\times 10^{-11}$. On the experimental side, the Muon $g-2$ Experiment at Fermilab released its Run-1 dataset in 2021 and the result was $a_{\mu}^{2021}(\mathrm{Exp})=116592061(41)\times 10^{-11}$ [5] when combining the data from Fermilab Run-1 and BNL. This resulted in a deviation from the SM prediction of $a_{\mu}^{2021}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})=(251\pm 59)\times 10^{-11}$ with a significance of $4.2\,\sigma$. In 2023, Fermilab released data from Run-2 and Run-3 leading to a new result after being combined with Run-1 and the BNL data as $a_{\mu}^{2023}(\mathrm{Exp})=116592059(22)\times 10^{-11}$ [6]. In this case the deviation from the SM prediction $\Delta a_{\mu}=a_{\mu}^{2023}(\mathrm{Exp})-a_{\mu}(\mathrm{SM})=(249\pm 48)\times 10^{-11}$ achieved a significance of $5.1\,\sigma$  [^1] Recently, there have been some discussions on the discrepancies in the calculation of the hadronic vacuum polarization contribution to $a_{\mu}(\mathrm{SM})$ between the lattice QCD calculations and experimental data in measurements of $e^{-}e^{+}\to\pi^{-}\pi^{+}$, which seems to reduce $\Delta a_{\mu}$ to some extend [7, 8, 9]. . These deviations inspired phenomenological investigations in various new physics models. Typical examples include the two-Higgs-doublet model [10, 11], the dark photon model [12], the supersymmetric models [13, 14, 15, 16, 17, 18, 14], leptoquark models [19, 20, 21], and vector-like lepton extended models [22, 23]. More references can be found in reviews [2, 3, 24, 25, 26, 27].
The chiral structure of the $a_{\mu}$ can be described by the following effective tensor operator
$\mathcal{L}^{\rm AMM}_{\rm eff.}=-\frac{e_{\ell}a_{\ell}}{4m_{\ell}}\,\overline{\ell_{L}}\,\sigma^{\mu\nu}\,\ell_{R}\,F_{\mu\nu}\,+{\rm h.c.}\,,$ (1)
@@
in which the SM charged lepton $\ell$ should be understood as muon flavor. We can see that both the left-handed (LH) $\ell_{L}$ and the right-handed (RH) $\ell_{R}$ of muon are involved. If one considers the new physics contribution to $a_{\mu}$ at 1-loop level matching the required chiral structure, proper chiral flip is needed along the fermion line. This chiral flip can take place either in the external muon line realized by the SM Yukawa interaction outside the loop or in the internal fermion line inside the loop [26].
@@
In the framework of simplified models, contributions to $\Delta a_{\mu}$ can be generated at the 1-loop level by introducing several new physics fields with various spins and quantum numbers under the electroweak gauge group ${\rm SU(2)_{L}\times U(1)_{Y}}$. As shown in [27], many models fail to explain $\Delta a_{\mu}$ either due to the wrong sign of predicted $\Delta a_{\mu}$ or excluded by other experimental constraints such as electroweak precision observables, flavor physics, and collider searches. Specifically, to our interest, it has been reported that simplified models introducing two new fields with different spins to the SM are unable to generate the chiral enhancement at the 1-loop level [27]. As a result, these models tend to confine the new physics particles to a low and compressed mass region in order to explain $\Delta a_{\mu}$ while avoiding collider constraints. However, an important consideration has been overlooked in the aforementioned analyses, i.e. the Yukawa interaction involving the SM Higgs doublet and the new physics fermion. The Yukawa interaction as a four-dimensional operator can naturally exist and induce mass mixing between the fermions in the SM and the new physics sector, thus providing an additional source of chiral flip to $a_{\mu}$.
@@
In this work, we demonstrate that by incorporating the SM Higgs Yukawa interaction into the analysis, simplified models introducing two new physics fields to the SM can effectively explain $\Delta a_{\mu}$. As a concrete example, we consider a simplified model that extends the SM with a scalar triplet $(3,-1)$ and a lepton triplet $(3,0)$ under the electroweak gauge group ${\rm SU(2)_{L}\times U(1)_{Y}}$. These two fields have been scrutinized in neutrino mass generation mechanisms as Type-II [28, 29, 30, 31, 32, 33] and Type-III [34, 35] seesaw models, respectively. However, concentrating specifically on the physics of $(g-2)_{\mu}$, in this work we do not require our model to produce the experimentally suggested texture of neutrino mass matrix. Instead, we only ensure the theoretically predicted neutrino masses remains to be negligibly small.
@@
This paper is organized as follows. In Sec. II we articulate our model setup and the related parameters. In Sec. III we present our analytical and numerical results. We draw our conclusion in Sec. IV.
@@ II.1 Model setup
In our model we extend the SM with a scalar triplet $S$ and a left-handed lepton triplet $F_{L}$, of which the representation $(n,\,Y)$ under the SM electroweak gauge group ${\rm\mathrm{SU(2)_{L}}\times\mathrm{U(1)_{Y}}}$ are
@@
$S\equiv\Big{[}\begin{array}[]{cc}\delta^{-}/\sqrt{2}&\quad(v_{\delta}+\delta^{0}+i\,a^{0})/\sqrt{2}\\ \delta^{--}&-\delta^{-}/\sqrt{2}\end{array}\Big{]}\,,\qquad F_{L}\equiv\Big{[}\begin{array}[]{cc}F_{L}^{0}/\sqrt{2}&F_{L}^{+}\\ F_{L}^{-}&-F_{L}^{0}/\sqrt{2}\end{array}\Big{]}\,,$ (7)
in which $v_{\delta}$ is the vacuum expectation value (vev) of the neutral component of $S$ after electroweak symmetry breaking (EWSB). We consider the following mass and Yukawa terms in the Lagrangian which are most relevant to the physics of $(g-2)_{\mu}$
@@
in which $H$ is the SM Higgs doublet and $\epsilon\equiv i\sigma^{2}$ is the antisymmetric tensor. $L^{i/j}_{L}\equiv(\nu^{i/j}_{L},\,\ell^{i/j}_{L})^{T}$ and $\ell^{i/j}_{R}$ are the LH doublet and RH singlet of SM lepton under ${\rm\mathrm{SU(2)_{L}}}$ with $i,j=1,2,3$ denoting the generation index, respectively. $(F_{L})^{C}$ denotes the charge conjugation of $F_{L}$ satisfying $(F_{L})^{C}\equiv C\overline{F_{L}}^{T}$. We require all Yukawa couplings in our model to be real in order to avoid constraints from CP-violating observables.
After EWSB we have the expansion $H=[G^{+},({v_{h}}+h+i\,G^{0})/\sqrt{2}]^{T}$ and can derive the following neutrino and charged lepton mass matrices
@@
Then the Yukawa interactions can be written as
@@
In the above expression, the replacement $\delta^{0}\to i\,a^{0}$ would generate the Yukawa interactions of $a^{0}$ and we do not write them explicitly for compactness. Moreover, as will be discussed later, $G^{0}$ ($G^{\pm}$) will mix with $\delta^{0}$ ($\delta^{\pm}$) and yield mass eigenstate with zero mass, corresponding to the Goldstone bosons absorbed to be the longitudinal components of SM gauge bosons $Z$ ($W^{\pm}$) via the Higgs mechanism. Therefore, we also do not write terms of $G^{0},G^{\pm}$ explicitly for compactness.
For simplicity, we consider the scenario that only the second generation $i,j\equiv 2$ exist in Eq. (8), i.e. the mass mixing in Eqs. (9) and (10) would generate mass eigenstates of charged lepton and neutrino in SM for muon flavor, as well as heavy neutral and charged leptons in the new physics sector. Therefore, the indices $i,j$ will be dropped in the following. We can diagonalize the previous mass matrices through the following rotations
@@
In the above, $s_{L/R}^{\nu/\ell}$ and $c_{L/R}^{\nu/\ell}$ are abbreviations of $\sin\theta_{L/R}^{\nu/\ell}$ and $\cos\theta_{L/R}^{\nu/\ell}$ when applicable. After the transformations in Eq. (11) we obtain the following Yukawa interactions in terms of mass eigenstates
@@
Again, note that the replacement $\delta^{0}\to i\,a^{0}$ in $\mathcal{L}_{\textrm{Yuk.}}$ would generate the Yukawa interactions of $a^{0}$ and we do not write them explicitly for compactness.
As for the scalar sector, there can be rich interactions involving $H$ and $S$ [36] as follows
$V(H,S)=$ $-m_{H}^{2}H^{{\dagger}}H+\frac{\lambda}{4}(H^{{\dagger}}H)^{2}+M_{S}^{2}\mathrm{Tr}[S^{{\dagger}}S]+[\mu_{HS}(H^{T}\epsilon SH)+\mathrm{h.c.}]$
@@
which would generate the mixing between $H$ and $S$ and result in the mass eigenstates including the SM Higgs boson and several scalars in the new physics sector. Similar to Eq. (11) we can diagnolize the electrically neutral CP-even scalar mass matrix by performing the following rotation
@@
$\left[\begin{array}[]{c}G^{0}\\ a^{0}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c^{a}&s^{a}\\ -s^{a}&c^{a}\end{array}\right]\left[\begin{array}[]{c}G^{0}\\ a^{0}\end{array}\right]\,,$
$\left[\begin{array}[]{c}(G^{+})^{\ast}\\ \delta^{-}\end{array}\right]$ $\rightarrow\left[\begin{array}[]{cc}c^{G}&s^{G}\\ -s^{G}&c^{G}\end{array}\right]\left[\begin{array}[]{c}(G^{+})^{\ast}\\ \delta^{-}\end{array}\right]\,,$
in which $s^{h/a/G},c^{h/a/G}$ are the abbreviations of $\sin\theta^{h/a/G},\cos\theta^{h/a/G}$ following the same convention as Eq. (11), and $G^{0},G^{\pm}$ are the Goldstone bosons which will be absorbed to be the longitudinal components of SM gauge bosons $Z,W^{\pm}$ via the Higgs mechanism. For later analysis we will label the masses of physical scalars in the mass eigenstates as $m_{h},m_{\delta^{0}},m_{a^{0}},m_{\delta^{-}},m_{\delta^{--}}$, with $h,\delta^{0}$ being the neutral CP-even scalars, $a^{0}$ being the neutral CP-odd scalar, and $\delta^{-},\delta^{--}$ being the singly and doubly charged scalars, respectively. Details of the diagnolization and rotation matrices in the scalar sector can be found in e.g. [36].
@@ II.2 Relations of parameters
Given that there are different equalities on $z_{L}v_{h}$ and $M_{F}$ in Eqs. (II.2) and (II.2), we can have the following identities
@@
  • $s_{R}^{\ell}$ characterizing the mixing of RH component of SM charged lepton $\ell$ with $F^{-}$ can be solved from Eqs. (22) and (II.2). In the parameter regions chosen for our numerical calculation it turns out to be the following  [^2] Note that in other models the suppression of $s^{\ell}_{R}$ compared to $s^{\ell}_{L}$ can be different from our model. Taking Type-III seesaw model as an example, the suppression is $\sim m_{\ell}/M_{F}$ [39, 40].
@@
    which means that the small angle approximation also holds for $\theta_{R}$ (see Eqs. (II.3) and (II.3) for more details). Therefore, in our model we have
@@
  • As the mass gap $\Delta m_{F}\equiv m_{F^{\pm}}-m_{F^{0}}$ in Eq. (24) is negligibly small, $\Delta m_{F}$ does not alter the main decay signals of the heavy leptons $F^{\pm},F^{0}$ as discussed in [41, 42]. Therefore, we can simply impose the similar lower bound of mass to be
@@
II.3 Input parameters
Now we determine the physically reasonable choice of input parameters of our model.
In the scalar sector shown in Eq. (II.1), despite the rich parameters and phenomenology about $V(H,S)$ (see e.g. [43] for collider signal searches), the physical quantities most relevant to $(g-2)_{\mu}$ are
$v_{\delta}\,,\,v_{h}\,,\qquad m_{h}\,,\,m_{\delta^{0}}\,,\,m_{a^{0}}\,,\,m_{\delta^{-}}\,,\,m_{\delta^{--}}\,,\qquad s^{h}\,,\,s^{a}\,,\,s^{G}\,,$ (30)
which are internally related [36]. Our requirements and simplifications include the following ones.
  • As for the two vevs $v_{h},v_{\delta}$, we require them to satisfy [44, 45]
    $v=\sqrt{v_{h}^{2}+2v_{\delta}^{2}}\approx 246\,{\rm GeV}\,,$ (31)
    to be consistent with the electroweak precision measurement [38] and
    $m_{h}=125\,{\rm GeV}\,,$ (32)
    to match the mass of the SM-like Higgs boson discovered at the Large Hadron Collider (LHC) [46, 47, 38]. To avoid the constraints $v_{\delta}\lesssim 5\,{\rm GeV}$ from the electroweak precision observables [48], in our numerical calculation we choose
    $v_{\delta}=5\,{\rm GeV}\,,\,v_{h}\approx 246\,{\rm GeV}\,.$ (33)
  • As for the masses of heavy scalars in the new physics sector, we require that the following approximations hold well
    $M_{S}\approx m_{\delta^{0}}\approx m_{a^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}\,,$ (34)
    which can be properly realized under the conditions of $v_{\delta}\ll v_{h}$ and $m_{h}\ll M_{S}$ [36]. Hereafter we would use $m_{\delta}$ as the notation to denote
    $m_{\delta}\equiv m_{\delta^{0}}\approx m_{a^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}\,.$ (35)
    Considering that the current searches for doubly charged scalar sets a lower limit of mass to be around $1\,{\rm TeV}$ assuming decaying to SM leptons [49, 50], we take the following benchmark throughout this work